OpenStudy (anonymous):

4-(6x^2-2)/(15)=(5x^2)/(3) - (5(x^2-2))/2

OpenStudy (anonymous):

\[\large4-\frac{6x^2-2}{15}=\frac{5x^2}{3}-\frac{5(x^2-2)}{2}\]??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i can't see any snappy way to do this, you might want to start with \[\frac{5x^2}{3}-\frac{5(x^2-2)}{2}+\frac{6x^2-2}{15}=4\] then multiply both sides by 30 otherwise you have to add the fractions

OpenStudy (anonymous):

yeah, that is probably best you get \[10\times 5x^2-15\times 5(x^2-2)+2(6x^2-2)=120\] then clean up the left and solve for \(x\)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

it really cleans up a lot you should end up i think with \[x^2=2\] but check that yw